alexgame:

The sum of two numbers is 47. The greater number is 9 more than the smaller number. Which equation can be used to solve for the smaller number? x + (x + 9) = 47 x − (x + 9) = 47 x(x + 9) = 47 x(x − 9) = 47

mhanifa:

Let the smaller number is x, then the greater number is...

mhanifa:

Then you add up the two numbers

mhanifa:

Can you do that?

alexgame:

hmmm

alexgame:

so you add 47 and 9

mhanifa:

47 is the sum of the two numbers

alexgame:

yes and

alexgame:

it could be any numbers

mhanifa:

If your smaller number is x then your greater number is 9 more than x, what would it be?

alexgame:

so the number has to be smaller than 9 right

SmokeyBrown:

@mhanifa wrote:
If your smaller number is x then your greater number is 9 more than x, what would it be?
(In terms of x, what is the bigger number? Remember, the greater number is 9 more than the smaller number, x)

alexgame:

i don't understand what you mean by greater number is 9 and smaller number x

mhanifa:

Let the smaller number is s and the greater number is g and their sum is 47. How would you write that?

SmokeyBrown:

mhanifa is saying that we can let "x" be the smaller number. Then, we know the bigger number is 9 more than x. So, how do we represent something being 9 more than x?

alexgame:

s + g =47

mhanifa:

@alexgame wrote:
s + g =47
Very good. Now, if g is 9 more than s, how would you write that?

alexgame:

9s + g = 47

alexgame:

ohh nooo

alexgame:

s + 9g = 47

mhanifa:

@alexgame wrote:
s + 9g = 47
9g means 9 times, but you need 9 more

mhanifa:

You should have: g = s + ...

alexgame:

g = s + 9

mhanifa:

@alexgame wrote:
g = s + 9
Excellent. So you have two numbers: s and s +9, their sum is 47 Consider changing s with x after, this will give you a required answer

alexgame:

so x + (x + 9) = 47

mhanifa:

@alexgame wrote:
so x + (x + 9) = 47
Congrats, you have the correct one.

alexgame:

yo thx for the seconde time

mhanifa:

No problem. I hope you have learned something

alexgame:

i sure did

alexgame:

sya